OpenStudy (anonymous):

Find the divergence of the vector field F(x,y) = (x^2 − y^2) i + 8xy j . div F =

OpenStudy (unklerhaukus):

\[\nabla\cdot \mathbf F(x,y) =\frac{\partial F_x}{\partial x}+\frac{\partial F_y}{\partial y}\]

OpenStudy (unklerhaukus):

\[F_x = x^2-y^2\implies\frac{\partial F_x}{\partial x} = ?\] \[F_y = 8xy\implies\frac{\partial F_y}{\partial y} = ??\]

OpenStudy (sidsiddhartha):

\[Div F=(F_x \hat{i}+F_y \hat{j}+F_z \hat{k})(\frac{ \partial }{ \partial x }\hat{i}+\frac{ \partial }{ \partial y }\hat{j}+\frac{ \partial }{ \partial z}\hat{k})\\\] that how the expression @UnkleRhaukus wrote is coming :)

OpenStudy (unklerhaukus):

@Ldaniel, can you take the partial derivatives of the components?

OpenStudy (sidsiddhartha):

\[\frac{ \partial }{ \partial x}=take~derivative~of~"x"~terms~keeping~"y"~as ~constant\]

OpenStudy (anonymous):

F_x= 2x F_y=8y

OpenStudy (sidsiddhartha):

yes correct :)

OpenStudy (unklerhaukus):

nope \[F_x = x^2-y^2\implies\frac{\partial F_x}{\partial x} = 2x\]\[F_y = 8xy\implies\frac{\partial F_y}{\partial y} =8x\]

OpenStudy (sidsiddhartha):

okey pit fall

OpenStudy (anonymous):

oh yeah sorry

OpenStudy (anonymous):

8xy(8x) +2x(x^2-y^2)?

OpenStudy (anonymous):

@sidsiddhartha

OpenStudy (anonymous):

so its 10x?

OpenStudy (unklerhaukus):

\[\nabla\cdot \mathbf F(x,y) \\ =\left(\frac{ \partial }{ \partial x }\hat{\boldsymbol\imath}+\frac{ \partial }{ \partial y }\hat{\boldsymbol\jmath}\right)\cdot\left(F_x \hat{\mathbf {\boldsymbol\imath} }+F_y \hat{\mathbf {\boldsymbol\jmath} }\right)\\ =\frac{\partial F_x}{\partial x}+\frac{\partial F_y}{\partial y}\] \[{\mathbf F}=(x^2 − y^2) \hat{\boldsymbol\imath} + 8xy \hat{\boldsymbol\jmath} \]\[F_x = x^2-y^2\implies\frac{\partial F_x}{\partial x} = 2x\]\[F_y = 8xy\implies\frac{\partial F_y}{\partial y} =8x\] \[\nabla\cdot \mathbf F(x,y) =2x+8x =10x\]